2.4
*Do not use journal entries format*
*There are 2 pictures A-G*


LO 2-1, 2-2 Effect of accounts receivable and accounts payable transactions on financial statements Exercise 2-4 The following events apply to Lewis and Harper, a public accounting firm, for the 2018 accounting period: 1. Performed $70,000 of services for clients on account 2. Performed $40,000 of services for cash. 3. Incurred $36,000 of other operating expenses on account.
76 Chapter 2 4. Paid $10,000 cash to an employee for salary. 5. Collected $47,000 cash from accounts receivable. 6. Paid $16,000 cash on accounts payable. 7. Paid an $8,000 cash dividend to the stockholders. 8. Accrued salaries were $2,000 at the end of 2016. Required a. Show the effects of the events on the financial statements using a horizontal statements model like the following one. In the Cash Flow column, use OA to designate operating activity, IA for inves- ment activity, FA for financing activity, and NC for net change in cash. Use NA to indicate the element is not affected by the event. The first event is recorded as an example. Assets Liabilities + Equity Cash Accts. Rec. Accts. Pay. Sal. Pay Ret. Earn. Rev. Exp.Net Inc. Cash Flow NA + 70,000NA = NA + NA + 70,00010,000 _NA_ 70,000 b. What is the amount of total assets at the end of 2018? c. What is the balance of accounts receivable at the end of 2018? d. What is the balance of accounts payable at the end of 2018? e. What is the difference between accounts receivable and accounts payable? f. What is net income for 2018? g. What is the amount of net cash flow from operating activities for 2018? NA
Answer #1

a) Effect of the events are shown as follows:- (Amounts in $)

Event No. Assets = Liabilities + Equity
Cash + Accts. Rec. = Accts. Pay. + Sal. Pay + Ret. Earn. Rev. - Exp. = Net Inc. Cash Flow
1) NA + 70,000 = NA + NA + 70,000 70,000 - NA = 70,000 NA
2) 40,000 + NA = NA + NA + 40,000 40,000 - NA = 40,000 40,000 OA
3) NA + NA = 36,000 + NA + -36,000 NA - 36,000 = -36,000 NA
4) -10,000 + NA = NA + NA + -10,000 NA - 10,000 = -10,000 -10,000 OA
5) 47,000 + -47,000 = NA + NA + NA NA - NA = NA 47,000 OA
6) -16,000 + NA = -16,000 + NA + NA NA - NA = NA -16,000 OA
7) -8,000 + NA = NA + NA + -8,000 NA - NA = NA -8,000 FA
8) NA + NA = NA + 2,000 + -2,000 NA - 2,000 = -2,000 NA
Total 53,000 + 23,000 = 20,000 + 2,000 + 54,000 110,000 - 48,000 = 62,000 53,000 NC

b) The amount of total assets at the end of 2018 = Cash + Accts. Rec.

= $53,000+$23,000 = $76,000

c) Accounts Receivable at the end of 2018 = $23,000

d) Accounts Payable at the end of 2018 = $20,000

e) Accounts Receivable - Accounts Payable = $23,000 - $20,000 = $3,000

f) Net Income for 2018 = $62,000

g) Net Cash flow from operating activities for 2018 = $40,000-$10,000+$47,000-$16,000

= $61,000